Green Grass, Inc. charges a flat rate of $15 per hour for lawn services. Their main competitor S & G Lawn
Care charges an initial fee of $30 plus $10 per hour. Find the break even point.

Answer by ikleyn(52780)

.
Green Grass, Inc. charges a flat rate of $15 per hour for lawn services. Their main competitor S & G Lawn
Care charges an initial fee of $30 plus $10 per hour. Find the break even point.
~~~~~~~~~~~~~~~

The equation to solve is THIS

    15*x = 30 + 10x,


where x denotes the hours till the break point.


From the equation

    15x - 10x = 30

       5x     = 30

        x     = 30/5 = 6 hours.    ANSWER

Solved.